OpenStudy (dinmosiaren):

Let = x2 – 16 and g(x) = x + 4. Find and its domain.

OpenStudy (dinmosiaren):

Your suppose to do fx/gx

OpenStudy (anonymous):

if f(x)=x^2 -16 ???

OpenStudy (dinmosiaren):

yeah

OpenStudy (anonymous):

f(x)/g(x)=[(x+4)(x-4)/(x+4)] =x-4 so the domain is \[(-\infty,\infty)\]

OpenStudy (dinmosiaren):

thanks

OpenStudy (dinmosiaren):

so is the domain all real numbers except \[x\neq4\] or \[x \neq-4\]

OpenStudy (dinmosiaren):

which one would be correct?

OpenStudy (dinmosiaren):

You there?

OpenStudy (anonymous):

when you ploting the graph y=x-4 so both of your graph is approach to infinity so. it can,t be \[x \neq 4\] so the value of domain is \[(-\infty,\infty)\]